For each of these items, just calculate the amount of depreciation expense for 2019. Each item below is separate case.

FORMULA: DEPRECIATION EXPENSE = COST x DEPRECIATION RATE

Adam brought a desk in 2016 for $2,000. He uses this desk 100% for his personal business. He decided to depreciate this item using the MACRS 150% decline HY convention

Alice purchased a race horse in 2018 for $5,000. She decided to depreciate this item using the MACRS 200% decline HY convention.
Answer #1

HY CONVENTION means the half year convention for deprecitaion is the deprecitaion schedule that treats all property acquired during the year as acquired exactly in the middle of the year.This means only half of the full year deprecition is allowed in the first year,while the remaining balance is deducted in the final year of the deprecition schedule ,or the year that property is sold

Desk is brought in 2016

the matematical of marcs depreciation is 7 years for desk

Depreciation for 2016=cost*1/useful life*A*half year convertion WHERE A=150%

=2000*(1/7)*150%*1/2=214

deprecition 2017=(cost-depreciation in previous year)*1/recoverable period*A=(2000-214)*(1/7)*150%=383

depreciation 2018=(2000-214-383)*(1/7)*150%=300

depreciation 2019=(2000-214-383-300)*(1/7)*150%=236.35

desk brought in 2018

the matematical of marcs depreciation is 3 years for race horses

Depreciation for 2018=cost*1/useful life*A*half year convertion where A=200%

=5000*1/3*200%*1/2=1666.66

deprecition 2019=(cost-depreciation in previous year)*1/recoverable period*A

=(5000-1666.66)*(1/3)*200%=2222.226